two car A and B are moving along stright road in the same direction with velocities of 25KM/HR and 40 KM/HR respectively find the velocity of car B relative to car A


1
Expert's answer
2021-02-22T10:20:21-0500
"v_{BA}=v_B-v_A\\v_{BA}=40-25=15\frac{km}{h}"